OpenStudy (anonymous):

|-17| + 5(2) = 20y - |-35| - 19y ^ THE 2 IS ON THE TOP RIGHT CORNER OF THE 5

OpenStudy (jessicap):

\[\left| -17 \right|+5^{2}=20y-\left| -35 \right|-19y\]Simplify the absolute values first. I used parentheses so that the negative signs don't get mixed up.\[(17)+5^{2}=2y-(35)-19y\]Next, simplify the exponents and combine like terms.\[17+25=2y-35-19y\]\[42=-35-17y\]Add 35 to both sides\[42+35=-35+35-17y\]\[77=-17y\]Divide by -17\[77\div-17=-17y \div-17\]\[-77/17=y\]Punch -77/17 into your calculator and get approx. -4.53.

OpenStudy (anonymous):

so the answer is -4.53 ? but that isnt any of my choices , my choices are A.-7 B.77 C.-77 D.7

OpenStudy (jessicap):

Oh sorry, I accidentally changed the 20 into a 2. If you rework it, you would get 20y-19y which equals 1y instead of-17y. This leaves you with y=77

OpenStudy (anonymous):

THANK YOU :)

OpenStudy (jessicap):

No problem! Glad to help.